# Detecting Discrepancies Between AI-Generated and Natural Images Using Uncertainty

### v1

@@ Abstract
In this work, we propose a novel approach for detecting AI-generated images by leveraging predictive uncertainty to mitigate misuse and associated risks. The motivation arises from the fundamental assumption regarding the distributional discrepancy between natural and AI-generated images. The feasibility of distinguishing natural images from AI-generated ones is grounded in the distribution discrepancy between them. Predictive uncertainty offers an effective approach for capturing distribution shifts, thereby providing insights into detecting AI-generated images. Namely, as the distribution shift between training and testing data increases, model performance typically degrades, often accompanied by increased predictive uncertainty. Therefore, we propose to employ predictive uncertainty to reflect the discrepancies between AI-generated and natural images. In this context, the challenge lies in ensuring that the model has been trained over sufficient natural images to avoid the risk of determining the distribution of natural images as that of generated images. We propose to leverage large-scale pre-trained models to calculate the uncertainty as the score for detecting AI-generated images. This leads to a simple yet effective method for detecting AI-generated images using large-scale vision models: images that induce high uncertainty are identified as AI-generated. Comprehensive experiments across multiple benchmarks demonstrate the effectiveness of our method.
@@ 1 INTRODUCTION
Recent advancements in generative models have revolutionized image generation, enabling the production of highly realistic images (Midjourney, ; Wukong, ; Rombach et al., 2022). Despite the remarkable capabilities of these models, they pose significant challenges, particularly the rise of deepfakes and manipulated content. The high degree of realism achievable by such technologies prompts urgent discussions about their potential misuse, especially in sensitive domains such as politics and economics. In response to these critical concerns, a variety of methodologies for detecting generated images have emerged. A prevalent strategy treats this detection task as a binary classification problem, necessitating the collection of extensive datasets comprising both natural and AI-generated images to train classifiers (Wang et al., 2020).
@@
While existing detection methods have demonstrated notable successes, they typically encounter challenges in generalizing to images produced by previously unseen generative models Wang et al. (2023a). One promising avenue to enhance the robustness of detection capabilities involves constructing more extensive training datasets by accumulating a diverse array of natural and synthetic images. However, these attempts are often computationally intensive, requiring substantial datasets for effective binary classification. Additionally, maintaining robust detection necessitates continually acquiring images generated by the latest models. And when the latest generative models are not open-sourced, acquiring a large number of generated images to train classifiers is challenging. This highlights the urgent need for a novel framework to detect AI-generated images without requiring AI-generated images.
@@
Figure 1: Large-scale fundamental models trained on a large number of real images are capable of distinguishing between real and generated images.
A recent work (Tan et al., 2024) shows that features extracted by ViT of CLIP (Radford et al., 2021) can be employed to separate natural and AI-generated images, motivating an effective approach to detecting images by training a binary classifier in the feature space of CLIP. This provides a promising direction to explore the possibility that large-scale foundational models already have the ability to capture the subtle differences between real images and AI-generated images. This may be similar to the emergence phenomenon in large language models (Wei et al., 2022), even if these models are not designed to distinguish between real and generated images. This phenomenon may arise from the training data of large vision models. In particular, since training models with AI-generated data usually leads to model collapse (Shumailov et al., 2024), current large vision models are mainly trained over natural data, making these models biased between natural and AI-generated data. Even though natural and AI-generated images appear extremely semantically similar to humans, such “biased” large models can capture the difference between natural and generated images.
The distribution discrepancy in features between natural and generated images motivates us to revisit the strategy of detecting AI-generated images. Specifically, the ability of humans to distinguish between natural and generated images relies on the existence of the discrepancy between them. Moreover, humans are not trained to distinguish natural and generated images. To verify this intuition, we visualize the features of natural and generated images using a large-scale visual fundamental model (DINOv2), following previous work (Tan et al., 2024). As shown in Figure 1, even for images sampled from the same class, there are large distributional discrepancies in the feature space of DINOv2.
@@
Building on this foundation, we propose to leverage predictive uncertainty as the score for distinguishing between natural and AI-generated images. This is because predictive uncertainty offers an effective approach for capturing distribution shifts. In particular, recent studies (Snoek et al., 2019; Schwaiger et al., 2020) indicate that models tend to show increased uncertainty for out-of-distribution (OOD) samples. Hence, for large vision models trained only on natural images, we can treat the natural images as in-distribution samples and the generated images as OOD samples. The challenge comes from efficiently obtaining the uncertainty of the model on the test samples. Classical approaches include Monte-Carlo Dropout (MC-Dropout) (Gal & Ghahramani, 2016) and Deep Ensembles (Lakshminarayanan et al., 2017). However, in our attempts, MC-Dropout obtains sub-optimal results (in Table 5), and it is not practical to train multiple large models independently for ensemble. Instead, we find that this uncertainty can be well captured by perturbing the weights of the models. As shown in Figure 2, when a moderate level of perturbation is applied, the real image has consistent features on the model before and after the perturbation, but the generated image has large differences in features on the model before and after the perturbation.
@@
In this paper, we propose a novel method for AI-generated image detection by weight perturbation (WePe). Our hypothesis is that the model has greater uncertainty in predicting the OOD sample compared to the ID sample, and that this uncertainty can be expressed through sensitivity to weight perturbations. For a large model trained on a large number of real images, the real images can be considered ID samples, while the generated images are considered OOD samples. Thus, the sensitivity of the samples to the weight perturbation of the large model can be an important indicator to determine whether the generative models generate the samples. Despite its simplicity, WePe achieves state-of-the-art performance on various benchmarks.
Figure 2: Natural and generated images are differently sensitive to perturbations in the model weights. When a moderate perturbation is applied ($0.1$), the natural image has essentially the same features on the model before and after the perturbation, but the generated image already shows a large difference.
@@
  • We provide a new perspective to detect AI-generated images by calculating predictive uncertainty. This is built upon an intuitive assumption that the natural and generated images differ in distribution, making it possible to employ uncertainty to represent the distribution discrepancy.
@@
  • We propose to leverage large vision models to calculate the predictive uncertainty. The intuition is that large vision models are merely trained on natural images, making it possible to exhibit different uncertainties about natural and generated images. We capture this uncertainty by weight perturbation to effectively detect images (Eq. 3).
  • Comprehensive experiments on multiple benchmarks demonstrate that the proposed method outperforms previous methods, including training-based methods, achieving state-of-the-art performance.
2 Related Works
AI-Generated images detection. Recent advancements in generative models, such as those by (Brock et al., 2019; Ho et al., 2020), have led to the creation of highly realistic images, highlighting the urgent need for effective algorithms to distinguish between natural and generated images. Prior research, including works by (Frank et al., 2020; Marra et al., 2018), primarily focuses on developing specialized binary classification neural networks to differentiate between natural and generated images. Notably, CNNspot (Wang et al., 2020) demonstrates that a standard image classifier trained on ProGAN can generalize across various architectures when combined with specific data augmentation techniques. F3Net (Qian et al., 2020) distinguishes real face images from fake face images with the help of frequency statistical differences. NPR (Tan et al., 2024) introduces the concept of neighboring pixel relationships to capture differences between natural and generated images. Although these methods show superior performance on generators in the training set, they often do not generalize well to unknown generators. In addition to this, training-based methods are susceptible to small perturbations in the image. For this reason, recently, some training-free methods have been proposed. AEROBLADE (Ricker et al., 2024) calculates the reconstruction error with the help of the autoencoder used in latent diffusion models (Rombach et al., 2022). RIGID (He et al., 2024) finds that real images are more robust to small noise perturbations than generated images in the representation space of the vision foundation models and exploits this property for detection. However, these methods usually make overly strong assumptions about natural or generated images, leading to insufficient generalization. In our paper, we propose a training-free detection method through uncertainty analysis. Based on the widespread phenomenon that generated images have greater uncertainty than real images on models trained with real images, our method achieves robust detection performance.
Uncertainty estimation. Uncertainty estimation in machine learning has seen significant advancements in recent years. (Gal & Ghahramani, 2016) introduces Monte Carlo Dropout (MC Dropout), which uses dropout at inference to estimate uncertainty from the variance of multiple predictions. (Lakshminarayanan et al., 2017) develops deep ensembles, demonstrating improved uncertainty estimates through training multiple model indepently with different initializations. Recent work by (Snoek et al., 2019) analyzes the calibration of uncertainty in deep learning models, highlighting the importance of reliable uncertainty measures. Additionally, (Guo et al., 2017) explore the use of temperature scaling to enhance the calibration of model predictions. In our paper, We measure uncertainty by perturbing the model’s weights.
Random weight perturbations. Weight perturbation, i.e. adding noise to network weights, has been extensively studied. Many methods (Khan et al., 2018; Wu et al., 2020) perform training regularization by adding noise to the weights of the neural network during training to improve the generalization of the network. Weight perturbation has also been applied in adversarial attacks to study the robustness of the network (He et al., 2019; Garg et al., 2020). In addition to this, some works also study the sensitivity of neural network to weights perturbation. (Cheney et al., 2017) shows that convolutional networks are surprisingly robust to a number of internal perturbations in the higher convolutional layers but the bottom convolutional layers are much more fragile. (Weng et al., 2020) proposes an efficient approach to compute a certified robustness bound of weight perturbations. In our work, we find neural networks trained on natural images exhibit different robustness to weight perturbations for natural and generated images. Based on this property, we propose an efficient algorithm to distinguish between natural and generated images.
3 Method
@@ 3.1 Motivation
Our method is built upon a foundational assumption that natural and generated images have different distributions. This is a reasonable assumption; otherwise, we cannot distinguish between natural and generated images. Fortunately, this assumption is consistent with previous work (Tan et al., 2024) and our empirical observations, as shown in Figure 1.
Hence, for the large models trained merely on natural images, we can regard natural images as in-distribution (ID) data while generated images as out-of-distribution (OOD) data. This distribution discrepancy can be reflected by the widely used predictive uncertainty, since neural networks typically exhibit higher uncertainty for OOD samples (Snoek et al., 2019; Schwaiger et al., 2020). This leads to a simple yet novel approach to determine whether a test image is generated by AI models when we can calculate its uncertainty on a pre-trained large vision model. In the following, we first give a complete introduction to our method WePe, then describe on our reasons for choosing the method, and finally, we explore why WePe works.
3.2 Uncertainty Definition
Classical methods of uncertainty estimation, such as Deep Ensembles and MC Dropout, can simply be viewed as using the variance of the results of multiple predictions as an estimation of uncertainty $u(\mathbf{x})$:
$\mu(\mathbf{x})=\frac{1}{n}\sum_{t=1}^{n}\hat{y}_{t}(\mathbf{x}),\quad u(\mathbf{x})=\sigma^{2}=\frac{1}{n}\sum_{t=1}^{n}\left(\hat{y}_{t}(\mathbf{x})-\mu(\mathbf{x})\right)^{2},$ (1)

### v2

# Epistemic Uncertainty for Generated Image Detection

@@ Abstract
We introduce a novel framework for AI-generated image detection through epistemic uncertainty, aiming to address critical security concerns in the era of generative models. Our key insight stems from the observation that distributional discrepancies between training and testing data manifest distinctively in the epistemic uncertainty space of machine learning models. In this context, the distribution shift between natural and generated images leads to elevated epistemic uncertainty in models trained on natural images when evaluating generated ones. Hence, we exploit this phenomenon by using epistemic uncertainty as a proxy for detecting generated images. This converts the challenge of generated image detection into the problem of uncertainty estimation, underscoring the generalization performance of the model used for uncertainty estimation. Fortunately, advanced large-scale vision models pre-trained on extensive natural images have shown excellent generalization performance for various scenarios. Thus, we utilize these pre-trained models to estimate the epistemic uncertainty of images and flag those with high uncertainty as generated. Extensive experiments demonstrate the efficacy of our method. Code is available at https://github.com/tmlr-group/WePe.
@@ 1 Introduction
Recent advancements in generative models have revolutionized image generation, enabling the production of highly realistic images (Midjourney,, 2022; Wukong,, 2022; Rombach et al.,, 2022). Despite the remarkable capabilities of these models, they pose significant challenges, particularly the rise of deepfakes and manipulated content. The high degree of realism achievable by such technologies prompts urgent discussions about their potential misuse, especially in sensitive domains such as politics and economics. In response to these critical concerns, a variety of methodologies for detecting generated images have emerged. A prevalent strategy treats this task as a binary classification problem, necessitating the collection of extensive datasets comprising both natural and AI-generated images to train classifiers (Wang et al.,, 2020).
@@
While existing methods have demonstrated notable successes, unseen generative models Wang et al., (2023) pose challenges for them in generalizing to images with distribution shifts. One promising avenue to enhance the robustness of detection capabilities involves constructing more extensive training sets by accumulating a diverse array of natural and generated images. However, these attempts are computationally intensive, requiring substantial datasets for effective classification. Besides, maintaining robust detection necessitates continually acquiring images generated by the latest generative models. And when the latest generative models are not open-sourced, acquiring a large number of generated images to train classifiers is challenging. This highlights the urgent need for a novel framework to detect generated images without reliance on generated images.
@@
Figure 1: Models trained on a large number of natural images are capable of distinguishing between natural and generated images.
A recent work (Tan et al.,, 2024) shows that features extracted by ViT of CLIP (Radford et al.,, 2021) can be employed to separate natural and AI-generated images, motivating an effective approach to detecting images by training a binary classifier in the feature space of CLIP. This provides a promising direction to explore the possibility that large-scale foundational models already have the ability to capture the subtle differences between natural images and AI-generated images. As shown in Figure 1, even for images sampled from the same class, there are large distributional discrepancies in the feature space of DINOv2. This observation is consistent with an important metric for evaluating generative models-the FID score. FID score measures feature distribution discrepancy between natural and generated images on the Inception network (Szegedy et al.,, 2015). A FID score of 0 indicates that there is no difference between the two distributions. However, even on these simple networks such as Inception v3, advanced generative models like ADM still achieve an FID score of 11.84, not to mention that on powerful models such as DINOv2, we observe significant feature distribution discrepancy. However, despite these distributional differences, the inherent diversity of natural images precludes direct modeling of their distribution. This challenge motivates us to consider an alternative approach to reflect the distributional differences between natural and generated images.
Figure 2: WePe reflects the distribution discrepancy between AI-generated and natural images.
@@
In this study, we exploit the distributional disparity between natural and generated images by leveraging epistemic uncertainty to differentiate them. Epistemic uncertainty quantifies a model’s confidence in its predictions, reflecting its knowledge of the data distribution. As the volume of training data increases, a model’s epistemic uncertainty for in-distribution (ID) data diminishes. For a foundational model pre-trained on an extensive dataset of natural images, we posit that its epistemic uncertainty is lower for natural images compared to generated ones, due to the model’s alignment with the natural distribution. This perception is consistent with recent studies (Snoek et al.,, 2019; Schwaiger et al.,, 2020), which indicate that models tend to show increased uncertainty for out-of-distribution (OOD) samples. The challenge comes from efficiently obtaining the uncertainty of the model on the test samples. Classical approaches include Monte-Carlo Dropout (MC-Dropout) (Gal and Ghahramani,, 2016) and Deep Ensembles (Lakshminarayanan et al.,, 2017). However, in our attempts, MC-Dropout obtains sub-optimal results (in Table 2), and it is challenging to train multiple large models independently for model ensemble. Instead, our theoretical results show that this uncertainty can be well captured by perturbing the weights of the models. As shown in Figure 3, when a moderate level of perturbation is applied, the natural image has consistent features on the model before and after the perturbation, but the generated image has large differences in features on the model before and after the perturbation.
@@
In this paper, we propose a novel method for AI-generated image detection by weight perturbation (WePe). Our hypothesis is that the model has greater uncertainty in predicting the OOD sample compared to the ID sample, and that this uncertainty can be expressed through sensitivity to weight perturbations. For a large model trained on a large number of natural images, the natural images can be considered ID samples, while the generated images are considered OOD samples. Thus, the sensitivity of the samples to the weight perturbation of the large model can be an important indicator to determine whether the sample is generated by the generative models or not. As shown in Figure 2, we calculate FID scores between different types of generated images and natural images on DINOv2. It shows the performance of WePe strongly correlates with FID score, indicating the effectiveness of WePe in detecting distribution discrepancy.
Figure 3: Natural and generated images exhibit distinct sensitivities to perturbations in model weights. A moderate perturbation ($0.1$) results in minimal changes to the features of the natural image, while the generated image shows significant differences.
@@
  • We provide a new perspective to detect AI-generated images by calculating predictive uncertainty. This is built on the fact that natural and generated images differ in data distribution, making it possible to employ uncertainty to represent the distribution discrepancy.
@@
  • Since the data distribution discrepancy between generated and natural images is reflected in feature distribution discrepancy on the vision model, we propose using a large vision model to compute prediction uncertainty to highlight this difference. The intuition is that large vision models are merely trained on natural images, making it possible to exhibit different uncertainties about natural and generated images. We capture this uncertainty through weight perturbation, enabling effective detection. (Eq. 13).
  • Extensive experiments across multiple benchmarks show the proposed method surpasses existing methods, highlighting the efficacy of uncertainty in detecting AI-generated images.
2 Preliminaries
2.1 AI-generated image detection
The task of AI-generated image detection involves classifying a test image $\mathbf{x}$ as either a natural image or one produced by a generative model. Traditional supervised learning approaches rely on a curated dataset comprising both natural and AI-generated images to train a feature extractor and a binary classifier. This process is formalized as follows.
Let $\mathcal{D}^{0}=\{\mathbf{x}^{0}_{1},\mathbf{x}^{0}_{2},\dots,\mathbf{x}^{0}_{N^{0}}\}$ denote a set of $N^{0}$ AI-generated images, each labeled as $y=0$ (generated), and $\mathcal{D}^{1}=\{\mathbf{x}^{1}_{1},\mathbf{x}^{1}_{2},\dots,\mathbf{x}^{1}_{N^{1}}\}$ denote a set of $N^{1}$ natural images, each labeled as $y=1$ (natural). The objective is to learn a feature extractor $F(\cdot;\theta_{F})$, parameterized by $\theta_{F}$, and a binary classifier $D(\cdot;\theta_{D})$, parameterized by $\theta_{D}$, by minimizing a loss function $\ell(\cdot)$ over the training data:
$F,D=\arg\min_{\theta_{F},\theta_{D}}\ell\left(D\left(F(\mathbf{x};\theta_{F});\theta_{D}\right),y\right),$ (1)
where $y\in\{0,1\}$ represents the ground-truth label of the input image $\mathbf{x}$.
Upon completion of training, the feature extractor and classifier are used to compute a decision score for a test image $\mathbf{x}$, defined as $S(\mathbf{x})=D(F(\mathbf{x};\theta_{F});\theta_{D})$. The source of the image is determined by comparing this score to a predefined threshold $\tau$:
$\operatorname{pred}(\mathbf{x})=\begin{cases}\text{generated},&\text{if }S(\mathbf{x})<\tau,\\ \text{natural},&\text{otherwise.}\end{cases}$ (2)
The robustness and generalization of such methods depend critically on the size and diversity of the training dataset. To address this, techniques such as CNNspot employ sophisticated data augmentation strategies, including Gaussian blur and JPEG compression, to enhance the variability of the training data. In contrast, UnivFD adopts a different approach by utilizing the CLIP model as a fixed feature extractor and training only a single linear classification layer. Despite these innovations, both methods exhibit limited generalization when applied to images produced by unseen generative models.
2.2 Bayesian neural networks and uncertainty estimation
Traditional neural networks map inputs $\mathbf{x}$ to outputs $\mathbf{y}$ through a parameterized function $f(\mathbf{x};\theta)$, where weights $\theta$ are optimized deterministically, offering no uncertainty quantification. Bayesian Neural Networks (Neal,, 2012) address this by modeling $\theta$ as random variables with a prior distribution $p(\theta)$. Given a dataset $\mathcal{D}$, the posterior is computed via Bayes’ theorem:
$p(\theta|\mathcal{D})=\frac{p(\mathcal{D}|\theta)p(\theta)}{p(\mathcal{D})}.$ (3)
For a new input $\mathbf{x}^{*}$, the posterior predictive distribution:
$p(\mathbf{y}^{*}|\mathbf{x}^{*},\mathcal{D})=\int\underbrace{p(\mathbf{y}^{*}|\mathbf{x}^{*},\theta)}_{\text{Aleatoric}}\underbrace{p(\theta|\mathcal{D})}_{\text{Epistemic}}d\theta$ (4)
captures prediction uncertainty.
Exact posterior inference is intractable due to the high dimensionality of $\theta$. Monte Carlo methods, such as MC Dropout (Gal and Ghahramani,, 2016), approximate the posterior by using dropout during both training and testing and generating multiple stochastic predictions. Uncertainty is quantified through the variance or entropy of these predictions, reflecting two types:
  • Epistemic Uncertainty: Uncertainty in model parameters, encoded in $p(\mathbf{w}|\mathcal{D})$, which decreases with more data.
  • Aleatoric Uncertainty: Inherent noise in the data, modeled by the likelihood $p(\mathbf{y}|\mathbf{x},\mathbf{w})$.
3 Uncertainty based AI-generated image detection
@@ 3.1 Motivation
Uncertainty estimation is a well-established field critical to deep learning practitioners, as it facilitates explicit handling of uncertain inputs and edge cases (Durasov et al.,, 2021; Everett et al.,, 2022). By quantifying model confidence, practitioners can make informed decisions, such as deferring high-uncertainty inputs to human evaluation in classification tasks, thereby improving model reliability and robustness. In this study, we do not propose a novel method for uncertainty estimation. Instead, we focus on distinguishing natural from generated images by leveraging epistemic uncertainty as a discriminative metric. Our approach is grounded in the well-established principle that (Lahlou et al.,, 2023; Gal and Ghahramani,, 2016):
Epistemic uncertainty, which reflects a model’s lack of knowledge, can be reduced by acquiring additional information.
Formally, we adopt a Bayesian framework to quantify epistemic uncertainty through the posterior distribution over model parameters:
$p(\theta|\mathcal{D})=\frac{p(\mathcal{D}|\theta)p(\theta)}{p(\mathcal{D})},$ (5)
where $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ denotes the training dataset, with samples drawn i.i.d. from $p(\mathbf{y}|\mathbf{x},\theta_{0})$. Epistemic uncertainty is captured by the posterior variance:
$\text{Var}(\theta|\mathcal{D})=\mathbb{E}_{\theta\sim p(\theta|\mathcal{D})}[(\theta-\mathbb{E}[\theta|\mathcal{D}])^{2}].$ (6)
Under regularity conditions, the Bernstein-von Mises theorem (Van der Vaart,, 2000) establishes that, as the sample size $N\to\infty$,
$p(\theta|\mathcal{D})\approx\mathcal{N}\left(\hat{\theta}_{N},\frac{1}{N}I(\theta_{0})^{-1}\right),$ (7)
where $\hat{\theta}_{N}$ is the maximum likelihood estimate, and $I(\theta_{0})=\mathbb{E}\left[\left(\frac{\partial\log p(\mathbf{y}|\mathbf{x},\theta)}{\partial\theta}\right)^{2}\big|_{\theta=\theta_{0}}\right]$ represents the Fisher information matrix. Here, $\theta_{0}$ approximates the true parameters of the natural image distribution, as learned by a pre-trained vision model. Consequently, $\text{Var}(\theta|\mathcal{D})\propto\frac{1}{N}$, indicating that epistemic uncertainty decreases as the training data volume increases.
For a test input $\mathbf{x}^{*}$ drawn from a distribution $p_{\text{test}}(\mathbf{x})\neq p_{\text{train}}(\mathbf{x})$, the predictive distribution is given by:
$p(\mathbf{y}^{*}|\mathbf{x}^{*},\mathcal{D})=\int p(\mathbf{y}^{*}|\mathbf{x}^{*},\theta)p(\theta|\mathcal{D})\,d\theta.$ (8)
This distribution exhibits elevated variance due to the misalignment between $p(\theta|\mathcal{D})$ and the parameters required to model $p_{\text{test}}(\mathbf{x})$. The predictive variance can be decomposed as:
$\text{Var}(\mathbf{y}^{*}|\mathbf{x}^{*},\mathcal{D})=\mathbb{E}_{\theta\sim p(\theta|\mathcal{D})}[\text{Var}(\mathbf{y}^{*}|\mathbf{x}^{*},\theta)]+\text{Var}_{\theta\sim p(\theta|\mathcal{D})}[\mathbb{E}(\mathbf{y}^{*}|\mathbf{x}^{*},\theta)],$ (9)
where the second term, $\text{Var}_{\theta\sim p(\theta|\mathcal{D})}[\mathbb{E}(\mathbf{y}^{*}|\mathbf{x}^{*},\theta)]$, corresponds to epistemic uncertainty and dominates for out-of-distribution (OOD) samples. Under distribution shift, the Fisher information $I(\theta_{0})$, defined with respect to $p_{\text{train}}(\mathbf{x})$, is misaligned with $p_{\text{test}}(\mathbf{x})$, limiting the reduction of posterior variance and resulting in persistently high epistemic uncertainty (Snoek et al.,, 2019).
Based on the above analysis, as well as the phenomenon of distributional shifts between natural and generated images on the foundational model (see Figures 1 and 8), we believe that the difference in the epistemic uncertainty of the foundational model on natural and generated images can be a valid indicator to distinguish between them.
3.2 Uncertainty estimation via weight perturbation
Classical methods of epistemic uncertainty estimation, such as Deep Ensembles and MC Dropout, can simply be viewed as using variance of multiple prediction results as an estimation of uncertainty $u(\mathbf{x})$:
$\mu(\mathbf{x})=\frac{1}{n}\sum_{t=1}^{n}\hat{y}_{t}(\mathbf{x}),u(\mathbf{x})=\sigma^{2}=\frac{1}{n}\sum_{t=1}^{n}\left(\hat{y}_{t}(\mathbf{x})-\mu(\mathbf{x})\right)^{2},$ (10)